# Any-point Trajectory Modeling for Policy Learning

### 28 Dec 2023 v1

@@ Abstract
Learning from demonstration is a powerful method for teaching robots new skills, and more demonstration data often improves policy learning. However, the high cost of collecting demonstration data is a significant bottleneck. Videos, as a rich data source, contain knowledge of behaviors, physics, and semantics, but extracting control-specific information from them is challenging due to the lack of action labels. In this work, we introduce a novel framework, Any-point Trajectory Modeling (ATM), that utilizes video demonstrations by pre-training a trajectory model to predict future trajectories of arbitrary points within a video frame. Once trained, these trajectories provide detailed control guidance, enabling the learning of robust visuomotor policies with minimal action-labeled data. Our method’s effectiveness is demonstrated across 130 simulation tasks, focusing on language-conditioned manipulation tasks. Visualizations and code are available at: https://xingyu-lin.github.io/atm.
@@
footnotetext: *First three authors contributed equally: Chuan Wen led the implementation and experiments. Xingyu Lin came up with the idea, supervised the technical development, and contributed to model debugging. John So implemented the first version of the code and the UniPi baselines.
1 Introduction
@@
Videos contain knowledge about behaviors, physics, and semantics, presenting an alternative data source. However, the lack of action labels makes utilization of video data in policy learning difficult. Previous works have addressed this by using self-supervised objectives for video pre-training to learn a feature representation of the observation for policy learning [39, 29, 37]. However, a feature representation only describes the state at the current time step, largely neglecting the transition dynamics that predicts future states. To explicitly model the transition dynamics, prior works have developed video prediction models that predict future image frames from current ones [11, 47, 12]. However, learning a video prediction model for control introduces two issues. Firstly, the task of video prediction avoids any abstraction by modeling changes to every pixel, coupling the physical motion with other factors such as texture, and lighting. This coupling makes modeling difficult, often resulting in hallucinations and unrealistic future predictions [11]. Secondly, these models are computationally demanding in both training and inferencing. With limited computational resources, performance significantly declines. Moreover, the high inference cost compels these models to adopt open-loop execution [11, 4], which tends to result in less robust policies.
@@
In this paper, we propose a novel and structured representation to bridge video pre-training and policy learning. We first represent each state as a set of points in a video frame. To model the temporal structure in videos, we learn an Any-point Trajectory Model (ATM) that takes the positions of the points in the current frame as input and outputs their future trajectories. We predict these trajectories in the camera coordinate frame to minimize the assumptions on calibrated cameras. These 2D point trajectories correspond to trajectories of particles in the 3D space and are a universal representation of the motions that can transfer to different domains and tasks. Contrasting with the video prediction approach of tracking changes in pixel intensity, our particle-based trajectory modeling offers a more faithful abstraction of the physical dynamics, naturally incorporating inductive biases like object permanence. We first pre-train the trajectory model on actionless video datasets. After pre-training, the predicted trajectories serve as detailed guidance for the policies, functioning like subgoals. We then train trajectory-guided policies using only a minimal amount of action-labeled demonstration data. For training the ATMs, we generate self-supervised training data by leveraging recent advancements in vision models for accurate point tracking [18]. We evaluate our method on a challenging simulation benchmark comprising 130 language-conditioned manipulation tasks [25]. Our experiments demonstrate that trajectory-guided policies significantly surpass various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. We summarize our main contributions below:
@@
    We propose the Any-point Trajectory Model, a simple and novel framework that bridges video pre-training to policy learning, leveraging the structured representation of particle trajectories.
@@
    Through extensive experiments on simulated benchmarks, we demonstrate that our method is able to effectively utilize video data in pre-training and significantly outperform various video pre-training baselines in an imitation learning setting.
2 Related Work
@@
State representation for control. In learning end-to-end visuomotor policies, the policy is typically parameterized as a neural network that takes image observation as the input representation [34, 5, 42]. Due to the lack of inductive bias, these approaches require training on a large number of demonstration trajectories, which is expensive to collect. On the other hand, different structured representations are proposed to improve the data efficiency, such as key points [32, 28], mesh [23, 17], or neural 3D representation [22, 33]. However, prior structures often limit the policy to specific tasks. In contrast, we propose to utilize future trajectories of arbitrary points in the image as additional input to the policy, providing an effective regularization. We demonstrate its wide application to a set of 130 language-conditioned manipulation tasks. In navigation and locomotion, it is common to construct policies that are guided by the future trajectories of the robot [1, 31]. In manipulation, some works have explored flow-based guidance [14, 36, 15]. However, prior works only track task-specific points, such as the center of mass of the robot or the end-effector of the robot. Instead, our approach works with arbitrary points in different camera viewpoints, thus providing richer information in the more general settings.
@@
Video pre-training for control. Videos contain rich information about behaviors and dynamics, which can help policy learning. However, video pre-training remains challenging due to the lack of action labels. One line of works first learns an inverse dynamics model that predicts the action from two adjacent frames and then labels the videos with pseudo actions [35, 2, 40]. Once labeled, the videos can be used for behavior cloning. However, the inverse dynamics model is often trained on a limited action-labeled dataset and does not generalize well, especially for continuous actions. The imprecise action labels will later hurt policy learning. Prior works have also explored pre-train a feature representation using various self-supervised objectives [38, 29, 26], but the representation alone does not retain the temporal information in videos. More recently, learning video prediction as pre-training has shown promising results [11, 21, 3, 47, 12]. During policy learning, a video prediction model is used to generate future subgoals and then a goal-conditioned policy can be learned to reach the sub-goal. However, video prediction models often result in hallucinations and unrealistic physical motions. Furthermore, video models require extensive computation, which is an issue, especially during inference time. In contrast, our method models the trajectories of the points, naturally incorporating inductive bias like object permanence while requiring much less computation. This enables our trajectory models to be run closed-loop during policy execution. Furthermore, the trajectories provide dense guidance to the policy as a motion prior.
Finally, Vecerik et al. [41] propose to utilize any-point tracking for few-short policy learning. This approach does not learn trajectory models from data, but instead mainly uses the tracker to perform visual servoing during test time. This design choice requires more instrumentation, such as separating the task into multiple stages, predicting the goal locations of the points for each stage, and running the tracker during inference time. In contrast, we present a much simpler framework, enabling application in more diverse settings.
@@ 3 Preliminary
Imitation from demos and videos. To begin with, we denote the action-free video dataset as $\mathcal{T}_{o}=\{(\tau_{o}^{(i)},\ell^{(i)})\}_{i=1}^{N_{o}},$ where $\ell^{(i)}$ is the language instruction for the $i^{th}$ episode and $\tau_{o}^{(i)}=\{o_{t}^{(i)}\}_{t=1}^{T}$ denotes the observation-only trajectory, mainly consist of camera images. Similarly, we denote the demonstration dataset as $\mathcal{T}_{a}=\{(\tau_{a}^{(i)},\ell^{(i)})\}_{i=1}^{N_{a}},$ where $\tau_{a}^{(i)}=\{o_{t}^{(i)},a_{t}^{(i)}\}_{t=1}^{T}$ is the action-labeled trajectory. During imitation learning, our goal is to learn a policy $\pi_{\theta}$, parameterized by $\theta$ to mimic the expert behavior by the following behavioral cloning objective:
@@ 4 Method
We propose Any-point Trajectory Modeling (ATM). As illustrated in Figure 2, ATM learns to predict future point trajectories in a video frame as the pre-training, then uses the predicted trajectories to guide policy learning. Our proposed method will be comprehensively detailed in this section: In Sec. 4.1, we first describe how to learn a point trajectory prediction model from an action-free video dataset $\mathcal{T}_{o}$. Then in Sec. 4.2, we outline how we utilize the pre-trained track prediction model to learn a track-guided policy from a limited action-labeled trajectory datasets $\mathcal{T}_{a}$.
4.1 Trajectory Modeling from Video Datasets
@@
Initially, we generate point trajectories from action-free videos for trajectory modeling pre-training. As described in Sec. 3, we employ a vision tracker to pre-process videos and generate a tracking dataset. For each video, we randomly sample a time step $\bar{t}$ and then randomly sample points on this frame and generate their tracks by running the tracker. However, for a static camera, most of the points that are sampled randomly will be in the background, thus providing little information when training our track transformer. To address this, we adopt a heuristic solution to filter out these static points: we first sample a grid of $n~{}\times~{}n$ on frame $\bar{t}$ and track the grid of points across the whole of the video to obtain an initial set of tracks $\tau\in\mathbb{R}^{n^{2}\times T\times 2}$. Subsequently, we filter points that have not moved during the video by thresholding the variance of the point positions over time. In the final step, we resample points around the filtered locations and finally generate their positions using the tracker.
@@
4.2 Track-guided Policy Learning
@@
While in training, we can filter out points without large movements to accelerate training, such a heuristic requires knowing the future positions of each point, which can be expensive to compute at test time. Instead, we find it sufficient to simply use a fixed set of 32 points on a grid for the policy. This sampling method avoids the potential complexities of learning key points or finding points to track [41] and works well in practice. ATM is permutation invariant to the input set of points, and we also find ATM to be robust to the distribution of the points, allowing us to use a different point sampling scheme from training for policy learning.
@@
5 Experiments
Figure 4: Task successes on the LIBERO benchmark. Our method significantly outperforms various baselines and approach the performance of BC trained on five times more demonstration data. We did not compare with UniPi on LIBERO-90 due to its high computation cost. The Overall success rates are computed over LIBERO-{Spatial, Object, Goal, Long}.
We conduct comprehensive experiments on 130 language-conditioned robotic manipulation tasks, to demonstrate the effectiveness of modeling point trajectories as video pre-training. Through extensive experiments, we aim to compare our method with strong baselines that pre-train from actionless videos. Additionally, we aim to demonstrate the importance of using a structured representation over RGB images. Finally, we aim to show the importance of various architectural design choices and provide visualizations to validate our claims.
5.1 Experiment Setup
@@
Environments. We compare with baselines on over one hundred language-conditioned manipulation tasks in the LIBERO benchmark [25]. The benchmark is subdivided into five suites, LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. Each suite focuses on different challenges in learning manipulation policies, such as the ability to reason about different scene layouts, object categories, task specifications, and long-horizon tasks. Each suite has 10 tasks, except LIBERO-90 which contains 90 tasks. Specifically, LIBERO-Spatial has tasks with the same objects but different layouts; LIBERO-Object has tasks with the same layouts but different objects; LIBERO-Goal has tasks with the same object categories and spatial layouts, but different goals; LIBERO-Long has tasks with diverse object categories and layouts, and long-horizon task goals; LIBERO-90 has extremely diverse object categories, layouts and task goals.
@@
Across all suites, each task contains $50$ demonstrations, including camera images as the observation. We only utilize the image observations for video pre-training. After video pre-training, all methods use 20% of the action-labeled demonstration trajectories (i.e., 10 demonstration trajectories for each task) to train behavioral cloning policies. The policies take RGB images from a third-person camera and a wrist camera, together with gripper and joint states as observations. As each task is specified by a language instruction, we use the pre-trained BERT network to obtain a task embedding [8]. In our experiments, the image resolution is $128\times 128$ and the action space is 7-dimension, representing the translation, rotation, and aperture of the end-effector.
Policy performance is measured by success rate: a score of 1 for complete task fulfillment, and 0 otherwise. Each policy is evaluated on 20 test rollout trajectories per task, and the success rates for each suite are averaged across all tasks. We report the mean and standard deviation of each suite’s average success rate across 3 seeds.
5.2 Baselines
Our work primarily concentrates on leveraging unlabeled video data to boost policy learning. In this context, we compare our method against four distinctive baselines for video pre-training:
@@ 5.2 Baselines
    BC denotes the vanilla behavioral cloning method which trains a policy exclusively using the limited action-labeled expert demonstrations. This method does not leverage the priors from videos.
@@
    VPT [2] first trains an inverse dynamic model from expert demonstrations on the video dataset $\mathcal{T}_{a}$ and then uses it to predict action labels for the video dataset. With these pseudo-labels, a policy is trained through behavioral cloning. This method requires the inverse dynamics model to be robust to a wide distribution of input observation, which can be difficult to learn from the limited demonstrations.
@@
    UniPi [11, 21, 4] trains a text-conditioned video diffusion model to generate a temporally fine-grained video plan from an initial frame and a language instruction. During policy learning, UniPi trains an inverse dynamic model with action-labeled data. We base our implementation off of the UniPi implementation in Ko et al. [21]. While both UniPi and ATM leverage a policy conditioned on future subgoals, a trajectory representation decouples motion from other pixel-based information and makes policy learning much easier. Please see the Appendix where we perform comprehensive comparisons on different variations of UniPi.
5.3 Results
@@
We present the main results in Figure 4. We see that by bridging the video data and policy learning with the structured representation of point trajectories, ATM (our method) significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$.
Compared with the vanilla BC baseline, ATM exhibits superior performance in all suites, showing significant knowledge transfer from the video dataset. R3M-finetune performs significantly worse than our method across all the suites, demonstrating the importance of explicitly modeling the transition dynamics for policy learning. As a representative paradigm for learning from videos, VPT exhibits the worst performance. We empirically observed that the pseudo-action labels predicted by VPT generally show higher errors than what can be learned from an end-to-end policy. Training with these imprecise action labels will further amplify the errors. We see that learning a good inverse dynamics model from limited demonstration data and for continuous action spaces can be challenging.
While the UniPi baseline performs similarly to ATM on the LIBERO-Spatial and LIBERO-Object, ATM significantly outperforms UniPi on LIBERO-Goal and LIBERO-Long, which requires understanding the tasks precisely and performing complex long-horizon tasks. Both UniPi and ATM share similar algorithmic components. While UniPi learns to propose future image observation as sub-goals, ATM uses future point trajectories as sub-goals. This comparison highlights the benefits of using a structured, trajectory sub-goal compared to an image goal.
From these comparisons, we found that ATM excels at the LIBERO-Spatial, LIBERO-Long, and LIBERO-90 suites. These tasks contain more diverse tasks and long-horizon tasks. ATM’s superior performance on these more challenging tasks shows its scalability to challenging robotics tasks and the importance of decoupling motions from pixel-based features.
5.4 Analysis
Next, we conduct a series of analysis experiments on LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long (omitting LIBERO-90 to save training and evaluation time) to explore the effect of the number of training demos and length of predicted tracks and ablate the components of track transformer and policy architectures.

### 16 Feb 2024 v2

@@ Abstract
[…]
@@
footnotetext: *First three authors contributed equally: Chuan Wen led the implementation and experiments. Xingyu Lin came up with the idea, supervised the technical development, and contributed to model debugging. John So implemented the Robot-to-robot transfer experiments and UniPi baselines.
I Introduction
@@
Videos contain knowledge about behaviors, physics, and semantics, presenting an alternative data source. However, the lack of action labels makes utilization of video data in policy learning difficult. Previous works have addressed this by using self-supervised objectives for video pre-training to learn a feature representation of the observation for policy learning [41, 31, 39]. However, a feature representation only describes the state at the current time step, largely neglecting the transition dynamics that predicts future states. To explicitly model the transition dynamics, prior works have developed video prediction models that predict future image frames from current ones [11, 51, 12]. However, learning a video prediction model for control introduces two issues. Firstly, the task of video prediction avoids any abstraction by modeling changes to every pixel, coupling the physical motion with visual appearances such as texture, and lighting. This coupling makes modeling difficult, often resulting in hallucinations and unrealistic future predictions [11]. Secondly, these models are computationally demanding in both training and inferencing. With limited computational resources, performance significantly declines. Moreover, the high inference cost compels these models to adopt open-loop execution [11, 5], which tends to result in less robust policies.
@@
In this paper, we propose a novel and structured representation to bridge video pre-training and policy learning. We first represent each state as a set of points in a video frame. To model the temporal structure in videos, we learn an Any-point Trajectory Model (ATM) that takes the positions of the points in the current frame as input and outputs their future trajectories. We predict these trajectories in the camera coordinate frame to minimize the assumptions on calibrated cameras. These 2D point trajectories correspond to trajectories of particles in the 3D space and are a universal representation of the motions that can transfer to different domains and tasks. Contrasting with the video prediction approach of tracking changes in pixel intensity, our particle-based trajectory modeling offers a more faithful abstraction of the physical dynamics, naturally incorporating inductive biases like object permanence. We first pre-train the trajectory model on actionless video datasets. After pre-training, the predicted trajectories serve as detailed guidance for the policies, functioning like subgoals. We then train trajectory-guided policies using only a minimal amount of action-labeled demonstration data. For training the ATMs, we generate self-supervised training data by leveraging recent advancements in vision models for accurate point tracking [18]. Across over 130 language-conditioned tasks we evaluated in both simulation and the real world, ATM significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. Additionally, we demonstrate that our method facilitates effective transfer learning from human videos and videos of a robot with a different morphology. We summarize our main contributions below:
@@
    We propose an Any-point Trajectory Model, a simple and novel framework that bridges video pre-training to policy learning, leveraging the structured representation of particle trajectories.
@@
    Through extensive experiments on simulated benchmarks and in the real world, we demonstrate that our method can effectively utilize video data in pre-training and significantly outperform various video pre-training baselines in an imitation learning setting.
  3. 3.
    We demonstrate effective learning from human videos.
II Related Work
@@
State representation for control. In learning end-to-end visuomotor policies, the policy is typically parameterized as a neural network that takes image observation as the input representation [36, 6, 46]. Due to the lack of inductive bias, these approaches require training on a large number of demonstration trajectories, which is expensive to collect. On the other hand, different structured representations are proposed to improve the data efficiency, such as key points [34, 29], mesh [23, 17], or neural 3D representation [22, 35]. However, prior structures often limit the policy to specific tasks. In contrast, we propose to utilize future trajectories of arbitrary points in the image as additional input to the policy, making minimal assumptions about the task and environment. We demonstrate its wide application to a set of over 130 language-conditioned manipulation tasks. In navigation and locomotion, it is common to construct policies that are guided by the future trajectories of the robot [1, 33]. In manipulation, some works have explored flow-based guidance [14, 38, 15]. However, prior works only track task-specific points, such as the end-effector of the robot, or the human hand. Instead, our approach works with arbitrary points, including points on the objects, thus providing richer information in the more general settings. Finally, Vecerik et al. [45] proposes to utilize any-point tracking for few-shot policy learning. This approach does not learn trajectory models from data, but instead mainly uses the tracker to perform visual servoing during test time. This design choice requires more instrumentation, such as separating the task into multiple stages, predicting the goal locations of the points for each stage, and running the tracker during inference time. In contrast, we present a much simpler framework, enabling application in more diverse settings.
@@
Video pre-training for control. Videos contain rich information about behaviors and dynamics, which can help policy learning. However, video pre-training remains challenging due to the lack of action labels. One line of works first learns an inverse dynamics model that predicts the action from two adjacent frames and then labels the videos with pseudo actions [37, 3, 44]. However, the inverse dynamics model is often trained on a limited action-labeled dataset and does not generalize well, especially for continuous actions. Prior works have also explored pre-train a feature representation using various self-supervised objectives [40, 31, 26], but the representation alone does not retain the temporal information in videos. More recently, learning video prediction as pre-training has shown promising results [11, 21, 4, 51, 12]. During policy learning, a video prediction model is used to generate future subgoals and then a goal-conditioned policy can be learned to reach the sub-goal. However, video prediction models often result in hallucinations and unrealistic physical motions. Furthermore, video models require extensive computation, which is an issue, especially during inference time. In contrast, our method models the trajectories of the points, naturally incorporating inductive bias like object permanence while requiring much less computation. This enables our trajectory models to be run closed-loop during policy execution. Furthermore, the trajectories provide dense guidance to the policy as a motion prior.
Learning from Human Videos. Of particular interest, numerous prior works have proposed learning from the rich source of human videos [50, 42, 30, 2, 43]. However, these works often explicitly extract the hand pose or contact regions from the human videos, thereby losing information about the dynamics of the remaining objects. In contrast, our method models the trajectories of arbitrary points and can learn from both human videos and videos of a different robot.
@@ III Preliminary
Imitation from demos and videos. To begin with, we denote the action-free video dataset as $\mathcal{T}_{o}=\{(\tau_{o}^{(i)},\ell^{(i)})\}_{i=1}^{N_{o}},$ where $\ell^{(i)}$ is the language instruction for the $i^{th}$ episode and $\tau_{o}^{(i)}=\{o_{t}^{(i)}\}_{t=1}^{T}$ denotes the observation-only trajectory consisting of camera images. Similarly, we denote the demonstration dataset as $\mathcal{T}_{a}=\{(\tau_{a}^{(i)},\ell^{(i)})\}_{i=1}^{N_{a}},$ where $\tau_{a}^{(i)}=\{o_{t}^{(i)},a_{t}^{(i)}\}_{t=1}^{T}$ is the action-labeled trajectory. During imitation learning, our goal is to learn a policy $\pi_{\theta}$, parameterized by $\theta$ to mimic the expert behavior by the following behavioral cloning objective:
@@ IV Method
We propose Any-point Trajectory Modeling (ATM). As illustrated in Figure 2, ATM learns to predict future point trajectories in a video frame as the pre-training, then uses the predicted trajectories to guide policy learning. Our proposed method will be comprehensively detailed in this section: In Sec. IV-A, we first describe how to learn a point trajectory prediction model from an action-free video dataset $\mathcal{T}_{o}$. Then in Sec. IV-B, we outline how we utilize the pre-trained track prediction model to learn a track-guided policy from a limited action-labeled trajectory datasets $\mathcal{T}_{a}$.
IV-A Trajectory Modeling from Video Datasets
@@
Initially, we generate point trajectories from action-free videos for trajectory modeling pre-training. As described in Sec. III, we employ a vision tracker to pre-process videos and generate a tracking dataset. For each video, we randomly sample a time step $\bar{t}$ and then randomly sample points on this frame and generate their tracks by running the tracker. However, for a static camera, most of the points that are sampled randomly will be in the background, thus providing little information when training our track transformer. To address this, we adopt a heuristic solution to filter out these static points: we first sample a grid of $n~{}\times~{}n$ on frame $\bar{t}$ and track the grid of points across the whole of the video to obtain an initial set of tracks $\tau\in\mathbb{R}^{n^{2}\times T\times 2}$. Subsequently, we filter points that have not moved during the video by thresholding the variance of the point positions over time. In the final step, we resample points around the filtered locations and finally generate their positions using the tracker.
@@
IV-B Track-guided Policy Learning
@@
During track transformer pre-training, we can filter tracks without large movements. However, using this heuristic requires knowing the future positions of each point, which can be expensive to compute during policy inference. Instead, we find it sufficient to simply use a fixed set of 32 points on a grid for the policy. This sampling method avoids the potential complexities of learning key points or finding points to track [45] and works well in practice. ATM is permutation invariant to the input set of points, and we also find ATM to be robust to the distribution of the points, allowing us to use a different point sampling scheme from training for policy learning.
@@
V Experiments
We perform experiments to answer the following questions:
  • How does ATM compare with state-of-the-art video pre-training baselines in learning from action-free videos?
  • Can ATM be used for learning from video data that are out of the distribution of demonstration data?
  • How do various design choices of ATM affect the performance?
Our experiments are split into three sections. In Sec. V-A, we compare ATM with video pre-training baselines on over 130 language-conditioned manipulation tasks in simulation and in the real world. In Sec. V-B, we show that ATM enables effective learning from human videos. Finally, we present ablation results in Sec. V-C.
Figure 4: We compare with state-of-the-art video pre-training methods on language-conditioned manipulation tasks in the LIBERO benchmark [25]. (a) Visualization of the LIBERO tasks separated into four suites, focusing on different aspects of the manipulation policies in spatial reasoning, object reasoning, task understanding, and performing long-horizon tasks. (b) Quantitative comparisons on different suites. We additionally compare baselines with fast computation on a task suite containing 90 tasks (i.e. LIBERO-90). ATM outperforms the baselines in all tasks and excels in LIBERO-Goal and LIBERO-Long.
Figure 5: Real robot experiments on a dining table setup consisting of five tasks. The left figure shows our real-world setup and the tasks. The top right figure shows an example of the predicted particle trajectories and the policy execution, which closely follows the predicted trajectories. From the quantitative results, we can see that ATM shows significant improvements over state-of-the-art video pre-training baselines on average.
V-A Video Pre-training for Imitation Learning
@@
Environments. We compare with baselines on over one hundred language-conditioned manipulation tasks in the LIBERO benchmark [25]. The benchmark is subdivided into five suites, LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. Each suite has 10 tasks, except LIBERO-90 which contains 90 tasks. Each task comes with expert human demonstrations. LIBERO-Spatial contains tasks with the same objects but different layouts; LIBERO-Object has tasks with the same layouts but different objects; LIBERO-Goal has tasks with the same object categories and spatial layouts, but different goals; LIBERO-Long has tasks with diverse object categories and layouts, and long-horizon task goals; LIBERO-90 has extremely diverse object categories, layouts and task goals.
@@
Data. We compare with baselines on each suite separately. All methods are trained on $10$ action-labeled demonstration trajectories and $50$ action-free video demonstration trajectories of the robot for each task, amounting to $500$ videos for each 10-task suite. The demonstration dataset contains RGB images from a third-person camera and a wrist camera, together with gripper and joint states as observations. As each task is specified by a language instruction, we use the pre-trained BERT network to obtain a task embedding [8]. Image resolution is $128\times 128$ and the action space is 7-dimension, representing the translation, rotation, and aperture of the end-effector.
Baselines. We compare with the following baselines:
@@
    BC denotes the vanilla behavioral cloning which trains a policy exclusively using the limited action-labeled expert demonstrations, without using the video dataset. It uses a policy architecture identical to ATM except that the particle trajectories are masked to be zero and it instead takes language embedding as task specification.
@@
    VPT [3] first trains an inverse dynamics model from the action-labeled trajectories $\mathcal{T}_{a}$ and then uses it to predict pseudo action labels for the video dataset. With these pseudo-labels, a policy is then trained with behavioral cloning. This method requires the inverse dynamics model to be robust to a wide distribution of input observation, which can be difficult to learn from the limited demonstrations.
@@
    UniPi [11, 21, 5] trains a text-conditioned video diffusion model to generate a temporally fine-grained video plan from an initial frame and a language instruction. During policy learning, UniPi trains an inverse dynamics model with action-labeled data. We base our implementation off of the UniPi implementation in Ko et al. [21]. While both UniPi and ATM leverage a policy conditioned on future subgoals, a trajectory representation decouples motion from other pixel-based information and makes policy learning much easier. Please see the Appendix where we perform additional comprehensive comparisons of different variations of UniPi.
@@
Results. We present the main results in Figure 4. We see that by bridging the video data and policy learning with the structured representation of point trajectories, ATM (our method) significantly surpasses various strong baselines in video pre-training, achieving an average success rate of $63\%$ compared to the highest success rate of $37\%$ by previous methods, marking an improvement of over $80\%$. The comparison of BC with ATM shows that learning from additional videos provides useful information for policy learning. VPT performs poorly as we empirically observed that the pseudo-action labels predicted by VPT generally show large errors on the video dataset. UniPi fails on more complex as video prediction models are not physically grounded and often generate future frames that are not physically feasible, such as cases where robots disappear from the image. Please see our video for failure cases of a video prediction model.
Real World Setup. We conduct a language-conditioned manipulation experiment in the real world to further strengthen our claim. As illustrated in Figure 5, we learn policies for a 6-DOF UR5 robot arm using human expert demonstrations collected with the GELLO teleoperation system [49]. The action space is joint position control and gripper state. The observations include two RGB images from two RealSense cameras. To compensate for the partial observation, we stack the most recent two frames as the input of agents. We do not feed proprioception states into agents because it is reported that imitation policies tend to overfit them and ignore the visual inputs, leading to worse online rollout performance [28, 24]. We collect a total of $50$ action-labeled trajectories and $250$ action-free demonstration videos (by simply removing the actions from the collected action-labeled demonstrations). From the quantitative results in Figure 5, we see consistent trends that ATM outperforms BC and other video pre-training baselines by a significant margin.
Figure 6: Learning robotic skills from human videos for three tasks. We collect 100 videos of a human performing the tasks directly and 10 teleoperation demonstration trajectories. Each row from the top to the bottom shows three snapshots from the human videos, ATM trained without the human videos, and ATM trained with the human videos. By comparison, we can see that human videos are critical in learning accurate trajectory prediction and enable the policy to successfully perform the task.
Figure 7: Cross-morphology skill transfer for a pick-and-place task. Here, we collect 160 action-free videos of a Franka arm and 10 action-labeled demonstrations from a UR arm, with the final goal of learning a UR policy. We compare a vanilla BC baseline with ATM trained using types of data: using only the 10 UR videos, using only the 160 Franka videos, and using both Franka and UR videos (Franka $\Rightarrow$ UR). In the right plot, we observe that the additional cross-embodiment data led to significantly better results. Surprisingly, even if the trajectory model is only trained using Franka videos, it exhibits much better performance than the BC without the Franka videos.
TABLE I: Average success rates of human-to-robot experiments. ATM trained with human videos significantly outperforms BC and ATM trained with only 10 robot videos, demonstrating the cross-embodiment capability of ATM.
Method
Teleoperation
demos
Human
videos
fold cloth put tomato sweep toys
BC 0% 10% 30%
ATM 0% 0% 13%
ATM 63% 63% 60%
V-B Human-to-robot and Robot-to-robot Transfer
By modeling the low-level any-point trajectories, ATM enables learning from cross-embodiment videos of humans or a different robot performing the task. This facilitates the use of more scalable data sources. To verify this, we present the results of learning from human videos in Fig. 6, with the quantitative results presented in Table I, and the results of cross-robot transfer in Fig. 7. We compare the following methods: (1) Learning behavior cloning policies only on the action-labeled data (2) ATM, using a track transformer trained only on the limited action-labeled robot data, and (3) ATM, using a track transformer trained on both the action-free human and action-labeled robot data. Experiments show that training the trajectory model on additional cross-embodiment videos makes the trajectory prediction more robust and accurate, significantly improving policy learning. On the other hand, as the number of action-labeled trajectories is small, BC baselines that only use action-labeled trajectories fail. Please refer to the videos on our website for better visualization.
V-C Ablation Analysis
We conduct a series of ablation experiments in simulation to demonstrate the effect of our design choices. These include the number of action-labeled trajectories needed for policy learning, the effect of the trajectory prediction horizon, image masking, and late track fusion.